1. Find the 3 consecutive odd integers such as the difference of 3x the third and 2x the second.
2. Ronnie is 30 yrs old and Grace is 22 yrs old. How many yrs ago was Ronnie twice as old as Grace.

Answer by macston(5194)

.
1.
This question is unclear (meaningless)
.
2.
R=Ronnie's age=30; G=Grace's age=22; x=years ago to condition
R-x=2(G-x)
30-x=2(22-x)
30-x=44-2x
x=14
.
ANSWER<:Fourteen years ago, Ronnie was twice as old as Grace.
.
CHECK:
R-14=2(G-14)
30-14=2(22-14)
16=2(8)
16=16
.